LUCI1 Commissioning

02:02 Tested new iif_binoc_runao command. Works

02:05 Tested pseudo-mono preset and then unchecked pseudo-mono button while the presets were executing. No problems and the telescope ended up in full binocular mode at the end of the preset.

02:45 LUCI binocular preset, seeing limited successful, after an issue with the script.

03:35 LUCI 1 preset works on 15 mag star, on second try. AO running in bin 4 @ 200Hz

04:32 Moved to 13 mag star. Bin 3. Optical loop data 20160425_043252

04:57 Removed actuator 557, DX, in current.

05:26 Closed for wind

-- %USERSIG{DougMiller - 2016-04-25}%
Topic revision: r3 - 25 Apr 2016, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback